How would you solve-2c+3(c+15)=c-7(-c+3)-9(-8+5c)?
Answer by mananth(16949)

-2c+3(c+15)=c-7(-c+3)-9(-8+5c)
-2c+3c+45=c+7c-21+72-45c
bring all the terms on LHS
-2c+3c-c-7c+45c=-45-21+72
38c=6
divide by 38
c= 6/38
c=3/19